To 250ml of M/50 sulphuric acid, 4g of solid sodium hydroxide is added and the resulting solution is diluted to 1litre. Determine the pH of the resulting solution

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

First of all determine teh equivalents of NaOH required to neutralise sulphuric acid and then we can get concentration of H+

take -logarithm of H+ to calculate pH of the resulting solution...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Devender Chopra
2 years ago
Moles of H2SO4 = n factor × Conc.×Volume
                                =  2×1/50×250/1000 
                                 =0.01 moles
Moles of NaOH = given mass / molar mass
                              =  4/40 = 0.1 moles 
Moles left of NaOH = 0.1-0.01 = 0.09 
                         [OH-] = 0.09 
pOH = -log(0.09) = 1.045
pH = 14-1.045 = 12.95 ← Ans.